OpenStudy (anonymous):

find the slope and y-intercept of the line write fractional answers in lowest term 5y+8x+2=5+8x

OpenStudy (anonymous):

could someone please help me find the slope and y-intercept of the line write fractional answers in lowest term 5y+8x+2=5+8x

OpenStudy (dumbcow):

get y by itself subtract 2 on both sides leaving 5y +8x = 3+8x subtract 8x on both sides leaving 5y = 3 divide by 5 leaving y = 3/5 since there is no x term, slope =0 y_intercept is then 3/5

OpenStudy (anonymous):

thanks so much dumbcow you made my day

OpenStudy (dumbcow):

:)